# Low-energy electron scattering from aminoacetonitrile using the R-matrix method

### v1

@@ 3.3 Excitation cross section
The electronic excitation cross sections of aminoacetonitrile into the lowest singlet and triplet electronic states were calculated using the $R$-matrix method with both the 6-311G* and cc-pVTZ basis sets, and the results are presented in figure 4. The graphs exhibit several distinct threshold behaviors and resonance-enhanced features between 7 and 10 eV. The excitation processes are symmetry-resolved in the upper panels and summed in the lower panels for the two basis sets.
@@
The cc-pVTZ basis yields significantly larger excitation cross sections and sharper resonance structures. The most prominent feature is a strong resonance in the $^{3}A^{\prime}$ channel, peaking near 8.5 eV with a maximum of $\sim 0.28\,\text{\AA }^{2}$. A second enhancement occurs at $\sim 9.5$ eV in both $^{3}A^{\prime}$ and $^{3}A^{\prime\prime}$ channels, producing a sharp peak in the total excitation cross section that exceeds $0.45\,\text{\AA }^{2}$. Compared with 6-311G*, the cc-pVTZ results clearly demonstrate that a more extensive basis set stabilizes excited states, lowers thresholds, and enhances coupling to the scattering electron, thereby producing stronger and more structured excitation signals.
The total excitation cross sections further emphasize the difference between the two basis sets. With 6-311G*, the overall magnitude is lower, and the resonance features appear as small bumps. In contrast, with cc-pVTZ, the total cross section rises more steeply above threshold, exhibits a pronounced peak at 8.5 eV, and shows an additional sharp maximum at 9.5 eV. These peaks align with the narrow high-energy resonances identified in the eigenphase analysis at 7–10 eV, supporting their interpretation as genuine, long-lived scattering states. This indicates that the more diffuse and flexible cc-pVTZ functions capture the coupling to excited states more accurately, leading to a clearer resonance signature. The presence of such strong features in the cc-pVTZ results also highlights the sensitivity of excitation channels to basis set quality, with 6-311G* underestimating resonance strengths and over-smoothing the excitation profile.
@@ 3.4 Momentum transfer cross section
The calculated momentum transfer cross sections (MTCS) for aminoacetonitrile are presented in figure 5 for the two basis sets, comparing the static-exchange (SE), static-exchange plus polarization (SEP), and configuration interaction (CI) models. At low energies, below 2 eV, all models yield very large MTCS values, consistent with strong forward scattering driven by the long-range dipole and polarization potential. As the energy increases, the MTCS values decrease sharply, reaching a pronounced minimum near 3 eV. Beyond this minimum, the MTCS exhibits resonance-driven enhancements in the 4–6 eV region. For the 6-311G* basis set, the SEP and CI models show clear peaks around 4.5–5 eV, while the SE model produces a broader rise centered near 6 eV. The cc-pVTZ results show a similar trend, with the SEP and CI models exhibiting peaks near 4.5–5 eV, reflecting the stabilization of the shape resonance identified in the eigenphase sum analysis. The SE model again produces a broader and less well-structured enhancement, consistent with its neglect of polarization effects.
At higher energies, from 7 to 10 eV, the MTCS values flatten to 110 - 130 $\text{\AA }^{2}$, with the CI models consistently predicting smaller cross sections than SE and SEP. The reduction in MTCS in CI models reflects more accurate cancellation between resonance contributions and background scattering, leading to weaker net momentum transfer. Additionally, the oscillatory behavior visible in the CI curves at higher energies corresponds to Feshbach resonances, reflecting coupling to inelastic channels that open in this energy region, in agreement with the excitation cross section results.
Overall, the MTCS results demonstrate the strong influence of target polarization and electron correlation effects. The SEP and CI models consistently reduce the magnitude relative to SE while introducing structure associated with low-lying resonances. These results, taken together with eigenphase, elastic, and excitation data, demonstrate that a careful treatment of correlation and basis-set convergence is essential for predicting momentum transfer processes, which are particularly relevant for modeling electron-driven transport and energy deposition in molecular environments such as interstellar ices or prebiotic gas mixtures.
@@ 3.5 Differential cross section
Figure 6 presents the differential cross sections (DCS) for elastic electron scattering from aminoacetonitrile, calculated at 1, 5, and 10 eV using the SE, SEP, and CI models with both basis sets (6-311G* and cc-pVTZ). At all energies the angular distributions are strongly forward-peaked, reflecting the dominance of long-range interactions at small scattering angles. The inclusion of polarization (SEP) and correlation (CI) generally lowers the DCS compared with SE, especially at intermediate and large scattering angles, consistent with enhanced absorption of flux into polarization and correlation channels.
@@
At 1 and 10 eV the three models show a very similar angular dependence, with only modest quantitative differences between SE, SEP, and CI. However, at 5 eV a distinct behavior emerges: while SEP and CI reduce the forward scattering relative to SE, at scattering angles beyond $90^{\circ}$ both SEP and CI exceed the SE result. This backward enhancement is not present at 1 or 10 eV, indicating it is tied to structure in the 4–6 eV energy region. The most plausible explanation is resonance-related: eigenphase analysis places low-lying resonances in this interval (between 4.5 and 5 eV), and inclusion of polarization and correlation modifies partial-wave phase shifts so that higher-order contributions interfere constructively at large angles. This behavior demonstrates the sensitivity of the angular distributions to resonance dynamics and the importance of polarization effects in describing backscattering.
Overall, the DCS results confirm the predominance of forward scattering in aminoacetonitrile, while also highlighting the role of resonance–polarization coupling in shaping the angular distributions at intermediate energies.
@@ 4 Conclusion
This work presents a detailed $R$-matrix study of electron scattering from aminoacetonitrile in the 0–10 eV range, reporting elastic, excitation, momentum transfer, and differential cross sections. By systematically comparing static-exchange (SE), static-exchange plus polarization (SEP), and configuration interaction (CI) models with two basis sets (6-311G* and cc-pVTZ), the role of polarization, correlation, and basis-set choice in shaping scattering dynamics has been established.
@@
The eigenphase analysis revealed a clear progression across models. SE calculations produced only diffuse, basis-dependent features without strong evidence of resonances. The inclusion of polarization in the SEP model stabilized and sharpened resonances, consistently locating them in the 4–5 eV region. The CI model provided the most reliable description, yielding robust $^{2}A^{\prime}$ and $^{2}A^{\prime\prime}$ resonances at $\sim$4.3–4.6 eV with widths of 1–1.7 eV, well reproduced across both basis sets. At higher energies (7–10 eV), extremely narrow resonances were detected in CI, suggesting possible long-lived Feshbach-type states.
@@
Elastic and momentum-transfer cross sections supported these resonance assignments. While SE produced smooth, featureless behavior, SEP and CI introduced structures near 4–5 eV that aligned with the low-energy eigenphase resonances. The momentum transfer cross sections further emphasized resonance-driven enhancements in the 4–6 eV region and showed overall reductions in magnitude when correlation was included. Excitation cross sections highlighted the importance of basis-set quality. With 6-311G*, excitation probabilities were underestimated, and resonance structures appeared smoothed out. In contrast, cc-pVTZ yielded significantly larger cross sections with pronounced peaks at 8.5 and 9.5 eV, consistent with the narrow high-energy resonances seen in eigenphase analysis. This underscores that diffuse basis functions are essential for accurately describing coupling to excited states. Differential cross sections confirmed forward-peaked scattering dominated by long-range interactions but also revealed resonance-driven modifications at 5 eV, particularly enhanced backscattering when polarization and correlation were included.
@@
In summary, this study demonstrates that reliable predictions of electron–aminoacetonitrile scattering require the explicit inclusion of polarization and correlation effects, supported by a sufficiently flexible basis set. The robust low-energy resonances at $\sim$4.5 eV are assigned as genuine scattering states and may facilitate temporary anion formation in environments where aminoacetonitrile is present, such as interstellar ices or prebiotic atmospheres. The additional high-energy resonances at 7–10 eV suggest further resonance-driven pathways for electronic excitation. The CI model with the cc-pVTZ basis set emerges as the most consistent and accurate framework, providing a solid foundation for future investigations into the role of electron attachment and excitation processes in the chemical evolution of prebiotic molecules.

### v2

@@ 3.3 Excitation cross section
The electronic excitation cross sections of aminoacetonitrile into the lowest singlet and triplet electronic states were calculated for both the 6-311G* and cc-pVTZ basis sets, and the results are presented in figure 4. The graphs exhibit several distinct threshold behaviours and resonance-enhanced features between 7 and 10 eV. The excitation processes are symmetry-resolved in the upper panels and summed in the lower panels for the two basis sets.
@@
The cc-pVTZ basis yields significantly larger excitation cross sections and sharper resonance structures. The most prominent feature is a strong resonance in the $^{3}A^{\prime}$ channel, peaking near 8.5 eV with a maximum of $\sim 0.28\,\text{\AA }^{2}$. A second enhancement occurs at $\sim 9.5$ eV in both $^{3}A^{\prime}$ and $^{3}A^{\prime\prime}$ channels, producing a sharp peak in the total excitation cross section that exceeds $0.45\,\text{\AA }^{2}$. Compared with 6-311G*, the cc-pVTZ results clearly demonstrate that a more extensive basis set stabilises excited states, lowers thresholds, and enhances coupling to the scattering electron, thereby producing stronger and more structured excitation signals.
The total excitation cross sections further emphasise the difference between the two basis sets. With 6-311G*, the overall magnitude is lower, and the resonance features appear as small bumps. In contrast, with cc-pVTZ, the total cross section rises more steeply above threshold, exhibits a pronounced peak at 8.5 eV, and shows an additional sharp maximum at 9.5 eV. These peaks align with the narrow high-energy resonances identified in the eigenphase analysis at 7–10 eV, supporting their interpretation as genuine, long-lived scattering states. This indicates that the more diffuse and flexible cc-pVTZ functions capture the coupling to excited states more accurately, leading to a clearer resonance signature. The presence of such strong features in the cc-pVTZ results also highlights the sensitivity of excitation channels to basis set quality, with 6-311G* underestimating resonance strengths and over-smoothing the excitation profile.
@@ 3.4 Momentum transfer cross section
The calculated momentum transfer cross sections (MTCS) for aminoacetonitrile are presented in figure 5 for the two basis sets, comparing the static-exchange (SE), static-exchange plus polarisation (SEP), and configuration interaction (CI) models. At low energies, below 2 eV, all models yield very large MTCS values, consistent with strong forward scattering driven by the long-range dipole and polarisation potential. As the energy increases, the MTCS values decrease sharply, reaching a pronounced minimum near 3 eV. Beyond this minimum, the MTCS exhibits resonance-driven enhancements in the 4–6 eV region. For the 6-311G* basis set, the SEP and CI models show clear peaks around 4.5–5 eV, while the SE model produces a broader rise centred near 6 eV. The cc-pVTZ results show a similar trend, with the SEP and CI models exhibiting peaks near 4.5–5 eV, reflecting the stabilisation of the shape resonance identified in the eigenphase sum analysis. The SE model again produces a broader and less well-structured enhancement, consistent with its neglect of polarisation effects.
At higher energies, from 7 to 10 eV, the MTCS values flatten to 110 - 130 $\text{\AA }^{2}$, with the CI models consistently predicting smaller cross sections than SE and SEP. The reduction in MTCS in CI models reflects more accurate cancellation between resonance contributions and background scattering, leading to weaker net momentum transfer. Additionally, the oscillatory behaviour visible in the CI curves at higher energies corresponds to Feshbach resonances, reflecting coupling to inelastic channels that open in this energy region, in agreement with the excitation cross section results.
Overall, the MTCS results demonstrate the strong influence of target polarisation and electron correlation effects. The SEP and CI models consistently reduce the magnitude relative to SE while introducing structure associated with low-lying resonances. These results, taken together with eigenphase, elastic, and excitation data, demonstrate that a careful treatment of correlation and basis-set convergence is essential for predicting momentum transfer processes, which are particularly relevant for modelling electron-driven transport and energy deposition in molecular environments such as interstellar ices or prebiotic gas mixtures.
@@ 3.5 Differential cross section
Figure 6 presents the differential cross sections (DCS) for elastic electron scattering from aminoacetonitrile, calculated at 1, 5, and 10 eV using the SE, SEP, and CI models with both basis sets (6-311G* and cc-pVTZ). At all energies the angular distributions are strongly forward-peaked, reflecting the dominance of long-range interactions at small scattering angles. The inclusion of polarisation (SEP) and correlation (CI) generally lowers the DCS compared with SE, especially at intermediate and large scattering angles, consistent with enhanced absorption of flux into polarisation and correlation channels.
@@
At 1 and 10 eV the three models show a very similar angular dependence, with only modest quantitative differences between SE, SEP, and CI. However, at 5 eV a distinct behaviour emerges: while SEP and CI reduce the forward scattering relative to SE, at scattering angles beyond $90^{\circ}$ both SEP and CI exceed the SE result. This backward enhancement is not present at 1 or 10 eV, indicating it is tied to structure in the 4–6 eV energy region. The most plausible explanation is resonance-related: eigenphase analysis places low-lying resonances in this interval (between 4.5 and 5 eV), and inclusion of polarisation and correlation modifies partial-wave phase shifts so that higher-order contributions interfere constructively at large angles. This behaviour demonstrates the sensitivity of the angular distributions to resonance dynamics and the importance of polarisation effects in describing backscattering.
Overall, the DCS results confirm the predominance of forward scattering in aminoacetonitrile, while also highlighting the role of resonance–polarisation coupling in shaping the angular distributions at intermediate energies.
@@ 4 Conclusion
This work presents a detailed $R$-matrix study of electron scattering from aminoacetonitrile in the $\sim$0–10 eV range, reporting integral elastic, excitation, momentum transfer, and differential elastic cross sections. By systematically comparing static-exchange (SE), static-exchange plus polarisation (SEP), and configuration interaction (CI) models with two basis sets (6-311G* and cc-pVTZ), the role of polarisation, correlation, and basis-set choice in shaping scattering dynamics has been established.
@@
The eigenphase analysis revealed a clear progression across models. SE calculations produced only diffuse, basis-dependent features without strong evidence of resonances. The inclusion of polarisation in the SEP model stabilised and sharpened resonances, consistently locating them in the 4–5 eV region. The CI model provided the most reliable description, yielding robust $^{2}A^{\prime}$ and $^{2}A^{\prime\prime}$ resonances at $\sim$4.3–4.6 eV with widths of 1–1.7 eV, well reproduced across both basis sets. At higher energies (7–10 eV), extremely narrow resonances were detected in CI, suggesting possible long-lived Feshbach-type states.
@@
Elastic and momentum-transfer cross sections supported these resonance assignments. While SE produced smooth, featureless behaviour, SEP and CI introduced structures near 4–5 eV that aligned with the low-energy eigenphase resonances. The momentum transfer cross sections further emphasised resonance-driven enhancements in the 4–6 eV region and showed overall reductions in magnitude when correlation was included. Excitation cross sections highlighted the importance of basis-set quality. With 6-311G*, excitation probabilities were underestimated, and resonance structures appeared smoothed out. In contrast, cc-pVTZ yielded significantly larger cross sections with pronounced peaks at 8.5 and 9.5 eV, consistent with the narrow high-energy resonances seen in eigenphase analysis. This underscores that diffuse basis functions are essential for accurately describing coupling to excited states. Differential cross sections confirmed forward-peaked scattering dominated by long-range interactions but also revealed resonance-driven modifications at 5 eV, particularly enhanced backscattering when polarisation and correlation were included.
@@
In summary, this study demonstrates that reliable predictions of electron–aminoacetonitrile scattering require the explicit inclusion of polarisation and correlation effects, supported by a sufficiently flexible basis set. The robust low-energy resonances at $\sim$4.5 eV are assigned as genuine scattering states and may facilitate temporary anion formation in environments where aminoacetonitrile is present, such as interstellar ices or prebiotic atmospheres. The additional high-energy resonances at 7–10 eV suggest further resonance-driven pathways for electronic excitation. The CI model with the cc-pVTZ basis set emerges as the most consistent and accurate framework, providing a solid foundation for future investigations into the role of electron attachment and excitation processes in the chemical evolution of prebiotic molecules.